The amount of money (y) that Juan has for any given week (x) can be modeled by the equation y = -7x + 80. About how long will it take for Juan to spend all of his money?
Answer by amarjeeth123(569)

The amount of money (y) that Juan has for any given week (x) can be modeled by the equation y = -7x + 80
y=0
-7x+80=0
x=80/7=11.4285 weeks
It will take 11.4285 weeks for Juan to spend all of his money.